(a) One litre of a sample of hard water contain 1 mg of CaCI2 and 1 mg of MgCI2. Find the total hardness in terms of parts of CaCO3 per 106 parts of water by weight.
(b) A sample of hard water contains 20 mg of Ca++ ions per litre. How many milli-equivalent of Na2CO3 would be required to soften 1 litre of the sample?
(c) 1 gm of Mg is burnt in a closed vessel which contains 0.5 gm of O2.
(i) Which reactant is left in excess?
(ii) Find the weight of the excess reactants.
(iii) How may milliliters of 0.5 N H2SO4 will dissolve the residue in the vessel.

Grade:10

4 Answers

Deepak Patra
9 years ago
. E = 12400 / λ (in Å) eV = 12400 / 975 = 12.75 eV … (i)
Also
13.6 [ 1 / n21 – 1 / n22] = 12. 75 ⇒ [1 /1 – 1 / n22] = 12.75 / 13. 6 ⇒ n2 = 4
For every possible transition one downward arrow is shown therefore the possibilities are 6.
Note : For longest wavelength, the frequency should be smallest.
This corresponds to the transition from n = 4 to n = 3, the energy will be E4 = 13.6 / 42 ; E3 = - 13.6 / 32
∴ E4 – E3 = 13.6 / 42 – (- 13.6 / 32) = 13.6 [1/9 – 1/16]
= 0.66 eV = 0.66 x 1.6 x 10-19 J = 1.056 x 10-19 J
Now, E = 12400 / λ (inÅ) eV ∴ λ = 18787 Å
Jitender Pal
9 years ago
(a) CaCI2 ≡ CaCO3 ≡ MgCI2
M. wt. 111 100 95
From this it is evident, the
111 mg CaCI2 will give CaCO3 = 100mg
∴ 1 mg CaCI2 will give CaCO3 = 100/111 mg = 0.90 mg
95 mg MgCI2 gives CaCO3 = 100mg
∴ 1 mg MgCI2 gives CaCO3 = 100/95 mg = 1.05 mg
∴ Total CaCO3 formed by 1 mg CaCI2 and 1 mg MgCI2 = 0.90 + 1.05 = 1.95 mg
∴ Amount of CaCO3 present per litre of water = 1.95mg
∴ wt of 1 ml of water = 1g = 103 mg
∴ wt of 100 ml of water = 103 * 103 = 106mg
∴ Total hardness of water in terms of parts of CaCO3 per 106 parts of water by weight = 1.95 parts.
(b) Eqwt of Ca++ = Mol.wt/Charge = 40/2 = 20
Ca2+ + Na2CO3 → CaCO3 + 2Na+
1 milliequivalent of Ca2+ = 20 mg
1 milliequivalentof Na2CO3 is required to soften 1 litre of hard water.
(c) 2Mg + O2 → 2MgO
2 * 24 = 48g 32g 2(24 + 16) = 80g
∵ 32g of O2 reacts with = 448g Mg
∴ 0.5g of O2 reacts with = 48/32 * 0.5 = 0.75g
Weight of unreacted Mg = 1.00 – 0.75 = 0.25g
Thus Mg is left in excess.
Weight of MgO formed = 80/48 * 0.75 = 1.25g
MgO + H2SO4 → MgSO4 + H2O
(40g)
According to reaction
∵ 40g MgO is dissolved it gives 1000 ml of 1 N. H2SO4
∴ 40 g MgO is dissolved it gives 2000 ml 0.5 N H2SO4
∴ 1.25 MgO is dissolved it gives
= 2000 *1.25/40 ml of 0.5 N H2SO4
= 62.5ml of 0.5N H2SO4
Ifjdndbishevehe
5 years ago
Hhhh should we just hang in the morning before you guys leave the school area in a little while I don’t want you going I have a lot to talk about tomorrow I have a good weekend thanks for your nice day and I hope you’re doing better thanks bye hello bye hi hello hi bye hello hello hi hello hello happy birthday Get low with the whistle blow 
Rishi Sharma
3 years ago
Dear Student,
Please find below the solution to your problem.

(a) CaCI2 ≡ CaCO3 ≡ MgCI2
M. wt. 111 100 95
From this it is evident, the
111 mg CaCI2 will give CaCO3 = 100mg
∴ 1 mg CaCI2 will give CaCO3 = 100/111 mg = 0.90 mg
95 mg MgCI2 gives CaCO3 = 100mg
∴ 1 mg MgCI2 gives CaCO3 = 100/95 mg = 1.05 mg
∴ Total CaCO3 formed by 1 mg CaCI2 and 1 mg MgCI2 = 0.90 + 1.05 = 1.95 mg
∴ Amount of CaCO3 present per litre of water = 1.95mg
∴ wt of 1 ml of water = 1g = 103 mg
∴ wt of 100 ml of water = 103 * 103 = 106mg
∴ Total hardness of water in terms of parts of CaCO3 per 106 parts of water by weight = 1.95 parts.
(b) Eqwt of Ca++ = Mol.wt/Charge = 40/2 = 20
Ca2+ + Na2CO3 → CaCO3 + 2Na+
1 milliequivalent of Ca2+ = 20 mg
1 milliequivalentof Na2CO3 is required to soften 1 litre of hard water.
(c) 2Mg + O2 → 2MgO
2 * 24 = 48g 32g 2(24 + 16) = 80g
∵ 32g of O2 reacts with = 448g Mg
∴ 0.5g of O2 reacts with = 48/32 * 0.5 = 0.75g
Weight of unreacted Mg = 1.00 – 0.75 = 0.25g
Thus Mg is left in excess.
Weight of MgO formed = 80/48 * 0.75 = 1.25g
MgO + H2SO4 → MgSO4 + H2O
(40g)
According to reaction
∵ 40g MgO is dissolved it gives 1000 ml of 1 N. H2SO4
∴ 40 g MgO is dissolved it gives 2000 ml 0.5 N H2SO4
∴ 1.25 MgO is dissolved it gives
= 2000 *1.25/40 ml of 0.5 N H2SO4
= 62.5ml of 0.5N H2SO4

Thanks and Regards